If I retire at age 62 I will have 28 years until then.

Based on the number of years until your retirement, calculate the monthly savings required to build a portfolio of $500,000. Assume you are starting with zero savings, can earn 6 percent a year, and that you start saving on the first of each month. What if you could earn 9 percent a year but wanted to save $750,000 – what would you need to deposit into savings monthly?
Answer #1

07 - e fix =PMT(C2,C3,C4,-C5, C6) в 0.005 0.06/12 336 28*12 ДА 1 2 RATE 3 NPER 4 PV 5 FV 6 TYPE 7 PMT Rate Per Period Total N

СРLUalu fx =PMT(C2,C3,C4,-C5,C6) B А 0.0075 0.09/12 336 28*12 1 2 RATE 3 NPER 4 PV 5 FV 6 TYPE 7 PMI Rate Per Period Total No